OpenStudy (roxy.girl):

Help me solve b + 4 = 7?

OpenStudy (roxy.girl):

@Qwertty123

rishavraj (rishavraj):

u tell wht u can do to isolate b

rishavraj (rishavraj):

i mean to to say get rid of "4" from the variable side ?? @Roxy.girl

OpenStudy (qwertty123):

Hi, This one isn't to hard... Reorder the terms: 4 + b = 7 Solving 4 + b = 7 Move all terms containing b to the left, all other terms to the right.

OpenStudy (roxy.girl):

You would add -4.

rishavraj (rishavraj):

good so u would add -4 on both sides..or u can say subtract 4 on both sides ...... so it would be b + 4 -4 = 7 -4

OpenStudy (qwertty123):

Then, Add '-4' to each side of the equation. 4 + -4 + b = 7 + -4 Combine like terms: 4 + -4 = 0 0 + b = 7 + -4 b = 7 + -4 Combine like terms: 7 + -4 =?

OpenStudy (roxy.girl):

oh okay.. so it would be 3?

rishavraj (rishavraj):

yup :))

OpenStudy (qwertty123):

Yes so b = 3!

OpenStudy (roxy.girl):

Thanks guys a few more! :)

OpenStudy (haseeb96):

You can ask me now.

OpenStudy (haseeb96):

What is your question @Roxy.girl